# Interference Among First-Price Pacing Equilibria: A Bias and Variance Analysis

### v1

@@
Abstract.
Online A/B testing is widely used in the internet industry to inform decisions on new feature roll-outs. For online marketplaces (such as advertising markets), standard approaches to A/B testing may lead to biased results when buyers operate under a budget constraint, as budget consumption in one arm of the experiment impacts performance of the other arm. To counteract this interference, one can use a budget-split design where the budget constraint operates on a per-arm basis and each arm receives an equal fraction of the budget, leading to “budget-controlled A/B testing.” Despite clear advantages of budget-controlled A/B testing, performance degrades when budget are split too small, limiting the overall throughput of such systems. In this paper, we propose a parallel budget-controlled A/B testing design where we use market segmentation to identify submarkets in the larger market, and we run parallel experiments on each submarket.
Our contributions are as follows: First, we introduce and demonstrate the effectiveness of the parallel budget-controlled A/B test design with submarkets in a large online marketplace environment. Second, we formally define market interference in first-price auction markets leveraging the first price pacing equilibrium (FPPE) framework which has been proposed to model autobidding in repeated auctions. Third, we propose a debiased surrogate that eliminates the first-order bias of FPPE, drawing upon the principles of sensitivity analysis in mathematical programs. Fourth, we derive a plug-in estimator for the surrogate and establish its asymptotic normality. Fifth, we provide an estimation procedure for submarket parallel budget-controlled A/B tests. Finally, we present numerical examples on semi-synthetic data, confirming that the debiasing technique achieves the desired coverage properties.
conference: ; ;
1. Introduction
Online A/B testing, or experimentation, is widely used in the internet industry. Major tech companies conduct a large number of A/B tests annually to inform decisions related to ads delivery, online auction mechanisms, user interface components, new feature rollouts, infrastructure improvements, and many other aspects of their products and services.
In the context of online auctions in marketplaces, where buyers bid for sellers’ items or advertisers bid for users’ impressions, budget-controlled A/B testing has gained popularity and is supported by companies like LinkedIn, Google, and Meta (Basse et al., 2016; Liu et al., 2021). The primary advantage of budget control in A/B testing is that it eliminates interference between the A and B groups that could otherwise bias the results: In a traditional A/B test, the budget of a buyer is distributed across all auctions, involving users from both A and B groups. This couples the behavior of the buyer across the groups; for example, increased expenditure under B might decrease expenditure in A, causing interference. However, in a budget-controlled A/B test, each buyer spends half of their budget on users from the A group and the other half on users from the B group, creating two separate but mirrored markets, and thus removing the interference caused by a shared budget constraint. This experiment type, therefore, is commonly used to test various ads delivery changes, such as pacing algorithms (see Section 2.1).
Despite clear advantages of budget-controlled A/B testing, companies are extremely constrained by the number of such experiments they can run: buyers (advertisers) can only be charged once for the same item (impression), it is thus not feasible to create additional experimentation throughput by reshuffling users between the A and B groups. An alternative idea is to create more budget splits so more treatments can be tested at the same time. However, this will lower the budget-per-group substantially, which could lead to different equilibrium outcomes and may disproportionately affect smaller buyers. In addition, it poses challenges with respect to statistical power, in view of the fact that tech companies primarily focus on down-funnel conversion events which are difficult to measure effects on to begin with. In fact, the lack of bandwidth for budget-controlled A/B tests has become one of the major bottlenecks in experimentation for the internet industry.
A third approach to increase throughput is market segmentation. If the overall market can be divided into several relatively isolated submarkets, budget-controlled A/B tests can be conducted in parallel within these submarkets. However, this method also presents some challenges. First, submarkets are rarely completely isolated; certain items may attract buyers from multiple submarkets, resulting in interference across submarkets when conducting tests in parallel. Second, submarkets differ in terms of buyer (and user) compositions, which might cause the local treatment effect estimates being not representative of the global treatment effect where all buyers are included in the market. The second challenge is relatively easy to address in practice by imposing balancing constraints in the clustering algorithm used to define submarkets, while the first challenge is more fundamentally important and requires deeper theoretical understanding.
As a motivating example, Fig 1(a) compares the results of submarket budget-controlled A/B tests (vertical axis) with those obtained from running budget-controlled A/B tests on the entire market (horizontal axis). The figure reports a normalized version of one of the real experiment metrics at Meta, with each submarket covering approximately 6-8% of advertisers and each experiment condition covering 30% of users. All submarkets are carefully created with limited imbalances in advertiser composition. We consider 16 different treatments, most of which are related to the pacing algorithm, each run for about 6 days, resulting in a total of 99 daily comparisons. The benchmark for sign-consistency at a 90% significance level (for neutral true effects) is 81.5%. The observed sign consistency rate between submarket and standard budget-controlled experiments is only slightly lower, at around 75%. Our focus on sign consistency in this context is intentional, as it directly relates to the frequency of production decision reversals, which we aim to minimize as much as possible.
(a) All datapoints.
(b) After guardrail metric thresholding. Note the different axis scale.
Figure 1. Submarket vs. standard budget-controlled A/B test, daily treatment effect. We denote neutral treatment effects with a value of $1.0$. For example, $1.01$ can be interpreted as $1\%$ treatment effect. Red crosses indicate instances of sign inconsistencies.
Table 1 shows the sign consistency rate of comparisons for daily and 6-day aggregated results, including for a second metric. Since submarkets are not perfectly isolated, we use a guardrail metric to measure the interference between submarkets, and use a threshold on the guardrail metric to improve reliability. As seen in Table 1, by eliminating experiments where the guardrail metric violates the threshold, we meaningfully improve the consistency rate for both metrics in daily and 6-day results. In practice, approximately 10% of submarket experiments violate the guardrail threshold. Fig 1(b) demonstrates the treatment effect comparisons for daily results after we remove such data points. The results are generally satisfactory, however the existence of biases is evident even after removing data points that violated the guardrail threshold. Evidently, quantitatively understanding the interference bias to improve the accuracy of the estimates is conducive to better decisions, and that is what we set to do in this work.
Table 1. Sign consistency between submarket and standard budget-controlled A/B test
Duration Data Num Sign consistency (%)
Metric 1 Metric 2
6 days All combos 16 87.5 68.8
6 days within guardrail 14 100 78.6
1 day All combos 99 75.8 73.7
1 day within guardrail 84 81.0 76.2
@@ 1. Introduction
The contributions of this paper are as follows. First, we introduce and demonstrate the effectiveness of the parallel budget-controlled A/B test design with submarkets in Meta’s production environment, which, to our best knowledge, is a novel approach. Second, we formally define market interference in first-price auction markets using the First Price Pacing Equilibrium (FPPE) framework (Conitzer et al., 2022a). Third, we propose a debiased surrogate that eliminates the first-order bias of FPPE, drawing upon the principles of sensitivity analysis in mathematical programs (Shapiro et al., 2021). Fourth, we derive a plug-in estimator for the surrogate and establish its asymptotic normality. Fifth, we provide an estimation procedure for submarket parallel budget-controlled A/B tests. Finally, we present numerical examples on semi-synthetic data, confirming that the debiasing technique achieves the desired coverage properties.
Section 2 introduces the FPPE framework. The formal problem setup is given in Section 3. Asymptotic theories around FPPE are presented in Section 4. Section 5 formalizes the parallel A/B testing problem under interference and provides the estimation procedure. Numerical examples on synthetic and semi-synthetic data are presented in Section 6. We conclude in Section 7.
1.1. Parallel A/B Testing in Practice
In this section we describe the real-world problem of A/B testing with congestion that we wish to model, and our proposed solution of parallel A/B tests in carefully balanced submarkets; the formal model appears in Sec. 2 and Sec. 3. Along with the description of the real-world problem, we will also discuss which aspects of it will be or will not be captured by our theoretical model.
@@ 1.1. Parallel A/B Testing in Practice
We start by describing the market environment. Assume that there is a set of $n$ advertisers, and each advertiser $i$ has a budget $b_{i}$. Whenever a user shows up on the platform an impression opportunity occurs, and an auction is conducted in order to determine which subset of ads will be shown to the user. Each advertiser $i$ has some stated value $v_{i}(\theta)$ of being shown to a particular impression opportunity $\theta$. This may not be their true valuation; in practice it may be their stated willingness to pay for a click, multiplied by an estimated click-through-rate for the pair ($i$,$\theta$). The advertiser submits a bid which is determined based on $v_{i}(\theta)$, as well as the expenditure of the advertiser so far. For example, in multiplicative pacing (Balseiro et al., 2017; Conitzer et al., 2022b, a), the platform adaptively learns a pacing multiplier $\beta_{i}\in[0,1]$ such that the bid is formed as $\beta_{i}v_{i}(\theta)$. The budget-management system then adaptively controls $\beta_{i}$ over time, in order to ensure the correct rate of budget expenditure on behalf of the advertiser. The auction format for an impression opportunity is typically some generalization of the first or second-price auction, though the real-world auctions are not fully captured by any simple-to-describe auction model from the literature. For example, in the setting of inserting ads into a news feed, multiple ads may be allowed to be shown, but the exact number may depend on the size of each ad, the size of other content in the feed, and so forth.
In our theoretical model, we will simplify the auction format in three respects: first, we will assume that each impression opportunity allows exactly one ad to be shown. This is a common assumption in the literature studying both online learning for budget management (Borgs et al., 2007; Balseiro and Gur, 2019) as well as in the literature on equilibrium outcomes under budget management (Balseiro et al., 2017; Conitzer et al., 2022b, a; Aggarwal et al., 2019; Balseiro et al., 2023). Thirdly, we assume that each individual auction is a first-price auction, meaning that the highest bidder wins, and pays their bid. Internet advertising uses both first and second-price auctions, depending on the particular context (Balseiro et al., 2023; Goke et al., 2021), though generalizations of the second-price auction are arguably still more common, in particular for sponsored search and feed auctions. We focus on first-price auctions because the equilibrium outcomes under budget management satisfy two crucial properties from the perspective of statistical inference: they are unique, and the equilibrium is captured by a convex program (Conitzer et al., 2022a). In contrast, equilibria under budget management in second-price auctions may not be unique (Conitzer et al., 2022b), and they are PPAD-hard to compute (Chen et al., 2023). It is an interesting future research direction to figure out how to develop results for the second-price setting, though we expect it to be much harder. We will also abstract away the dynamic temporal aspects of budget management: instead of the pacing multipliers being learned over time, we will assume that what we observe is the set of pacing multipliers in equilibrium. Secondly, in the real world, advertisers may use a variety of budget management algorithms beyond multiplicative pacing. For example, Google also offers probabilistic throttling, where bids are not modified, and instead the expenditure is managed by independently flipping a coin for each auction, in order to determine whether to bid at all (Balseiro et al., 2017; Chen et al., 2021; Karande et al., 2013). We assume that every advertiser uses multiplicative pacing in order to manage budgets. Budget-management equilibria in first-price auctions with probabilistic throttling are known to also be unique and computable (Chen et al., 2021), so it is possible that similar results could be obtained for that setting, though it is not known whether those equilibria are captured by a convex mathematical program, which is a crucial fact for our technical approach.
Now that we have discussed budget management, we describe the A/B testing problem. Suppose that a platform wants to run $K$ A/B tests, which may affect, e.g., the valuations that advertisers have for impression slots, revenue, etc. Naively performing the A/B test on a per-impression or per-user basis and then simply running the market as usual will not work, because the budget management system means that the A and B users will be affected by each other through the budget coupling. As mentioned in the introduction, a possible option is to perform budget splitting: create $K+1$ copies of each advertiser (corresponding to each test, as well as a holdout “A” control copy), randomize users into one of the $K+1$ markets, and run separate budget management for each copy of the $K+1$ advertisers. As mentioned earlier, this approach runs into problems because small advertisers end up with prohibitively-small budgets, and there is a large negative impact on statistical power because some of the events measured by the A/B tests might be far downstream of an impression.
@@
Instead, our paper considers the following parallel A/B test paradigm based on market segmentation. We first define a bipartite graph between advertisers and users based on targeting criteria. Subsequently, we cluster the advertisers into $K$ clusters, where the objective is to minimize the sum of weighted edges between clusters subject to traffic balancing constraints to make the resulting clusters as similar to the whole market as possible. The edge weight between a pair of advertisers is the number of impressions (or users) where they are both within the top-$k$ bids. The choice of $k$ is a parameter that must be chosen based on experience with the specific application setting. If the clustering achieves a small objective function value, then each cluster is a mostly isolated submarket, in the sense that each user will mostly receive bids from advertisers in a single cluster. Then, we run an A/B test within each of the $K$ submarkets. Every user is randomly assigned to either “A” or “B” in each submarket. Note that the same user can be in “A” in submarket 1 and “B” in submarket 2. This means that all the A/B tests must be mutually compatible; in practice this means that our approach is best suited for A/B tests that apply to advertisers, as opposed to, e.g., user-side user interface changes. If the interaction between different submarkets is not too strong, then we expect that hypothesis testing in a submarket is indicative of the true treatment effect within that submarket. Secondly, if each submarket is representative of the overall market, then such a hypothesis test is also indicative of the true treatment effect across the whole market.
In our theoretical model we use a coarser notion of contamination between a pair of submarkets than the top-$k$ measure that we use in practice. In our theoretical model, we measure the number of items with any bids from buyers from separate submarkets. In practice, this measure is too coarse (and often very large). It is an interesting follow-up question to come up with stronger theoretical results that allow for finer-grained measures of contamination. Secondly, our theoretical model will assume that what we are interested in is the treatment effect within a submarket. This implicitly assumes that each submarket is representative of the overall market.
2. Theory overview
2.1. Related Work
Pacing equilibrium. Pacing is a budget management strategy where bids are scaled down by a multiplicative factor in order to smooth out and control spending. In the first-price setting, Borgs et al. (2007) study first price auctions with budget constraints in a perturbed model, where the limit prices converge to those of an FPPE. Building on that work, Conitzer et al. (2022a) introduce the FPPE framework to model autobidding in repeated auctions, and discover several properties of FPPE such as shill-proofness monotonicity properties, and a close relationship between FPPE and the quasilinear Fisher market equilibrium (Chen et al., 2007; Cole et al., 2017). Gao and Kroer (2022) propose an infinite-dimensional variant of the quasilinear Fisher market, which lays the probability foundations of the current paper. Gao et al. (2021) and Liao et al. (2022) study online computation of the infinite-dimensional Fisher market equilibrium, and we utilize their methods to compute equilibria. In the second-price setting, a variety of models have been proposed for modeling the outcome of budget management: Balseiro et al. (2015) study budget-management in second-price auctions through a fluid mean-field approximation; Balseiro and Gur (2019) investigate adaptive pacing strategy from buyers’ perspective in a stochastic continuous setting; and Balseiro et al. (2021b) study several budget smoothing methods including multiplicative pacing in a stochastic context. A second-price analogue of FPPE was explored by Conitzer et al. (2022b). Modeling statistical inference in the second-price setting is most likely harder, due to equilibrium multiplicity issues (Conitzer et al., 2022b), and hardness of even computing equilibria (Chen et al., 2023).
Recently, a series of papers (e.g., Aggarwal et al. (2019); Deng et al. (2021); Balseiro et al. (2021a)) investigated an “autobidding” model for value-maximizing bidders, who aim to maximize their value subject to budget or return-on-spend constraints, which may include buying individual ad impressions that have negative utility. Several extensions provide results for more general settings, such as the quality of equilibria for generalized rationality models for bidders (Babaioff et al., 2021), non-uniform bidding strategies (Deng et al., 2023), and liquid welfare guarantees when budget-constrained buyers fail to reach equilibrium (Gaitonde et al., 2023). Statistical inference for parallel A/B testing under such models represent an interesting future direction of work.
Interference in A/B Testing for Auction Markets. Experimentation under interference has been extensively studied in the past decade within the context of social networks (Ugander et al., 2013; Aronow and Samii, 2017; Eckles et al., 2017; Athey et al., 2018; Li and Wager, 2022) and two-sided marketplaces (Zigler and Papadogeorgou, 2018; Wager and Xu, 2021). Under interference, the outcome of an experiment unit, either a user in a social network or a buyer in an online marketplace, can be affected by the treatment status of other units in the networks, violating the Stable Unit Treatment Values Assumption (SUTVA) assumption commonly employed in analyzing online A/B tests (Rosenbaum, 2007). To take into account this type of interference, novel designs and analysis approaches have been proposed, e.g., cluster randomization (Ugander et al., 2013; Karrer et al., 2021), regression-based estimators (Chin, 2019), and bipartite analysis (Harshaw et al., 2021). A/B testing in the presence of seller-side budget constraints has been discussed by Basse et al. (2016) and Liu et al. (2021) where the budget-splitting design spans the full market. However, most of the above works do not consider the particular type of interference caused by competitive equilibrium effects. Recent work has also addressed equilibrium effects (Wager and Xu, 2021; Liao et al., 2023; Liao and Kroer, 2024). Most relevant to our results, Liao and Kroer (2023) consider A/B testing via budget splitting in Fisher markets and FPPE. However, none of these works consider the effects caused by parallel A/B tests that interfere with each other. Our method builds on top of the budget-splitting design; our approach additionally segments the full market into submarkets, where different submarkets receive different treatment. In the formal sections of our paper we assume that the submarket structure is known. In practice, the submarkets may not be known; Rolnick et al. (2019) and Viviano et al. (2023) present methods to identify submarkets through clustering approaches that aim to limit the interference between clusters.
2.2. Review of FPPE
@@
Given a vector $a=[a_{1},\dots,a_{n}]^{\scriptscriptstyle\mathsf{T}}$, let ${\operatorname*{Diag}}(a)$ denote the diagonal matrix with $(i,i)$-th entry being $a_{i}$; sometimes we write ${\operatorname*{Diag}}(a_{i})$ when it is convenient to define each $a_{i}$ inline. Let $A^{\dagger}$ denote the Moore–Penrose inverse of the matrix $A$, $e_{j}$ the $j$-th unit vector, and $[n]=\{1,\ldots,n\}$.
Figure 2. Finite FPPE (left) and limit FPPE (right). In a finite FPPE, there are a finite number of items. In a limit FPPE, the item set is a continuum.
2.2.1. Basic Definitions, Limit FPPE and Regularity Conditions
We introduce the definitions of limit FPPE and two regularity conditions on the market. We have $n$ buyers and a possibly continuous set of items $\Theta$ with an integrating measure $\mathop{}\!\mathrm{d}\theta$. For example, one could take $\Theta=[0,1]$ and $\mathop{}\!\mathrm{d}\theta=$ the Lebesgue measure on $[0,1]$. An FPPE requires the following elements.
  • The budget $b_{i}$ of buyer $i$. Let $b=(b_{1},\dots,b_{n})$.
  • The valuation for buyer $i$ is a function $v_{i}\in L^{1}_{+}$, i.e., buyer $i$ has valuation $v_{i}(\theta)$ (value per unit supply) of item $\theta\in\Theta$. Let $v:\Theta\to{\mathbb{R}^{n}}$, $v(\theta)=[v_{1}(\theta),\dots,v_{n}(\theta)]$. We assume ${\bar{v}}=\max_{i}\sup_{\theta}v_{i}(\theta)<\infty$.
@@ 2.2.1. Basic Definitions, Limit FPPE and Regularity Conditions
  • The supplies of items are given by a function $s\in L^{\infty}_{+}$, i.e., item $\theta\in\Theta$ has $s(\theta)$ unit of supply. Without loss of generality, we assume a unit total supply $\int_{\Theta}s\mathop{}\!\mathrm{d}\theta=1$. Given $g:\Theta\to{\mathbb{R}}$, we let ${\mathbb{E}}[g]=\int g(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta$ and ${\operatorname{Var}}[g]={\mathbb{E}}[g^{2}]-({\mathbb{E}}[g])^{2}$. Given $t$ i.i.d. draws $\{\theta^{1},\dots,\theta^{t}\}$ from $s$, let $P_{t}g({\cdot})=\frac{1}{t}{\sum_{\tau=1}^{t}}g({\theta^{\tau}})$.
@@
Next we introduce the market equilibrium concepts that are the foundation of our study. For that we leverage the first-price pacing equilibrium (FPPE) concept put forward by Conitzer et al. (2022a). FPPE models equilibrium outcomes under budget-management systems employed in several practical settings. Each buyer is assigned a pacing multiplier $\beta_{i}$, which is used to control their budget expenditure. For each individual auction $\theta$, the buyer bids $\beta_{i}v_{i}(\theta)$, which can be seen as their adjusted valuation after factoring in their budget constraint (in practice, the valuation $v_{i}(\theta)$ may not be the buyer’s true valuation, but instead their unscaled bid reported to the platform). The goal of the budget management system is to achieve an equilibrium, meaning that it must ensure that buyers spend their budget exactly by appropriately choosing $\beta_{i}$. If the budget cannot be fully spent, then no pacing must occur (i.e., $\beta_{i}=1$).
We now formally define the pacing equilibrium concepts in the case of infinite (below) and finite number of users (in the next section), as illustrated by Fig 2.
Definition 1 (Limit FPPE, Gao and Kroer (2022); Conitzer et al. (2022a)).
@@
A limit FPPE, denoted as ${\small\mathsf{{FPPE}}}(b,v,s,\Theta)$, is the unique tuple $(\beta,p({\cdot}))\in[0,1]^{n}\times L^{1}_{+}(\Theta)$ such that there exist $x_{i}:\Theta\to[0,1]$ for all $i\in[n]$,
  1. (1)
@@
    (First-price) Prices are determined by first price auctions: for all item $\theta\in\Theta$, $p(\theta)=\max_{i}{\beta_{i}}v_{i}(\theta)$. Only highest bidders obtain items: for all $i$ and $\theta$, $x_{i}(\theta)>0$ implies ${\beta_{i}}{v_{i}(\theta)}=\max_{k}\beta_{k}v_{k}(\theta)$
  2. (2)
@@
    (Feasiblity, market clearing) Buyers are budget constrained: for all $i$, $\int x_{i}(\theta)p(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta\leq b_{i}$. There is no overselling: for all $\theta$, ${\sum_{i=1}^{n}}x_{i}(\theta)\leq 1$. All items are for which there is competition are fully allocated: for all $\theta$, $p(\theta)>0$ implies ${\sum_{i=1}^{n}}x_{i}(\theta)=1$.
  3. (3)

### v3

@@
Abstract
@@
A/B testing is widely used in the internet industry. For online marketplaces (such as advertising markets), standard approaches to A/B testing may lead to biased results when buyers have budget constraints, as budget consumption in one arm of the experiment impacts performance of the other arm. This is often addressed using a budget-split design. Yet such splitting may degrade statistical performance as budgets become too small in each arm. We propose a parallel budget-controlled A/B testing design where we use market segmentation to identify submarkets in the larger market, and we run parallel budget-split experiments in each submarket. We demonstrate the effectiveness of this approach on real experiments on advertising markets at Meta. Then, we formally study interference that derives from such experimental designs, using the first-price pacing equilibrium framework as our model of market equilibration. We propose a debiased surrogate that eliminates the first-order bias of FPPE, and derive a plug-in estimator for the surrogate and establish its asymptotic normality. We then provide an estimation procedure for submarket parallel budget-controlled A/B tests. Finally, we present numerical examples on semi-synthetic data, confirming that the debiasing technique achieves the desired coverage properties.
1 Introduction
Online A/B testing is widely used in the internet industry to inform decisions on new feature roll-outs. For online marketplaces (such as advertising markets), standard approaches to A/B testing may lead to biased results when buyers operate under a budget constraint, as budget consumption in one arm of the experiment impacts performance of the other arm. To counteract this interference, one can use a budget-split design where the budget constraint operates on a per-arm basis and each arm receives an equal fraction of the budget, leading to “budget-controlled A/B testing,” see e.g. Basse et al. (2016); Liu et al. (2021).
Despite clear advantages of budget-controlled A/B testing, companies are extremely constrained by the number of such experiments they can run. While it’s possible to create more budget splits, this will lower the budget per group substantially, which could lead to different equilibrium outcomes and may disproportionately affect smaller buyers. Additionally, a common approach to increase experimentation throughput is to run orthogonal experiments (with their own orthogonal randomization), but this would either suffer from the same interference as the vanilla A/B test setup, or also require further budget splits.
In this paper, we propose a parallel budget-controlled A/B test design where we use market segmentation to identify submarkets in the larger market, and we run parallel experiments on each submarket. When the overall market can be divided into several relatively isolated submarkets, budget-controlled A/B tests can be conducted in parallel within these submarkets. However, this method also presents some challenges. First, submarkets are rarely completely isolated; certain items may attract buyers from multiple submarkets, resulting in interference across submarkets when conducting tests in parallel. Second, submarkets differ in terms of buyer (and user) composition, which might cause the local treatment effect estimates to not be representative of the global treatment effect where all buyers are included in the market. The second challenge is relatively easy to address in practice by imposing balancing constraints in the clustering algorithm used to define submarkets, while the first challenge is more fundamentally important and requires deeper understanding.
Before the theoretical exposition, we consider a comparison of results for paired experiments between a parallel budget-controlled A/B test setup, and that of a traditional budget-split design; where the latter is considered the gold standard. Fig 1 on the left shows comparisons of 99 experiments where the point estimate and CIs are plotted on the vertical axis for the parallel design, and on the horizontal axis for the budget-split design. The most important feature is whether the two experiments agree between (negative, neutral, positive), as a change would result in a launch reversal. The two experiment designs agree in $75\%$ of cases (at $90\%$ confidence level, hence the optimal agreement is $81.5\%$), which increases to $79\%$ after the introducing a guardrail metric, see Fig 1 on the right. These results are quite satisfactory, but do point at the existence of remaining interference bias. In the remainder of this paper, correcting the interference bias is the main objective.
Figure 1: Parallel vs. standard budget-controlled A/B test, daily treatment effect. We denote neutral treatment effects with a value of $1.0$. Red crosses indicate instances of sign inconsistencies. Left are all datapoints, on the right, datapoints that fail a guardrail metric are removed.
Contributions:
  1. 1.
    We formally define market interference in first-price auction markets using the first price pacing equilibrium (FPPE) framework (Conitzer et al., 2022a) (Sec. 3)
  2. 2.
    We propose a debiased surrogate that eliminates the first-order bias of FPPE, and derive a plug-in estimator for the surrogate and establish its asymptotic normality. (Sec. 4)
  3. 3.
    We run semi-synthetic experiments, confirming that the debiasing technique achieves the desired coverage properties. (Sec 5).
1.1 Parallel A/B Testing in Practice
@@
In this section we describe the real-world problem of A/B testing with congestion that we wish to model, and our proposed solution of parallel A/B tests in carefully balanced submarkets. We start by describing the market environment. There is a set of $n$ advertisers, and each advertiser $i$ has a budget $b_{i}$. Whenever a user shows up on the platform an impression opportunity occurs, and an auction is conducted in order to determine which ad will be shown to the user. Each advertiser $i$ has some stated value $v_{i}(\theta)$ of being shown to a particular impression opportunity $\theta$. The advertiser submits a bid which is determined based on $v_{i}(\theta)$, as well as the expenditure of the advertiser so far. For example, in multiplicative pacing Balseiro et al. (2017); Conitzer et al. (2022b), the platform adaptively learns a pacing multiplier $\beta_{i}\in[0,1]$ such that the bid is formed as $\beta_{i}v_{i}(\theta)$. The budget-management system then adaptively controls $\beta_{i}$ over time, in order to ensure the correct rate of budget expenditure on behalf of the advertiser. We consider first-price auctions, which is the predominant way display advertising is sold online.
@@
Now that we have discussed budget management, we describe the A/B testing problem. Suppose that a platform wants to run $K$ A/B tests, which may affect, e.g., the valuations that advertisers have for impression slots, revenue, etc. We construct the market-segmented experimental setup as follows: We first define a bipartite graph between advertisers and users based on targeting criteria. Subsequently, we cluster the advertisers into $K$ clusters, where the objective is to minimize the sum of weighted edges between clusters subject to traffic balancing constraints to make the resulting clusters as similar to the whole market as possible. The edge weight between a pair of advertisers is the number of impressions (or users) where they are both within the top-$k$ bids. The choice of $k$ is a parameter that must be chosen based on experience with the specific application setting. If the clustering achieves a small objective function value, then each cluster is a mostly isolated submarket, in the sense that each user will mostly receive bids from advertisers in a single cluster. Then, we run an A/B test within each of the $K$ submarkets. Every user is randomly assigned to either “A” or “B” in each submarket. The main challenge is that while submarkets are relatively isolated, there is remaining interference from users who are targeted by advertisers from different submarkets, leading to a slightly different equilibrium. Our main contribution is to define a framework for analyzing such interference, and giving an estimator that removes the bias from these users. We survey related works in App. D.
2 Review of FPPE Theory
@@
Figure 2: Left: Finite FPPE (left) and limit FPPE (right). In a finite FPPE, there are a finite number of items; in a limit FPPE, the item set is a continuum. Right: The interference model — Left (^$\widehat{\mathscr{M}}_{\alpha}$): the observed market where interference is present among submarkets. Middle (${\mathscr{M}}_{\alpha}$): the limit market with interference from bad item set. Right (${\mathscr{M}}_{0}$): the limit market with perfectly separated submarkets. We use data from the left panel to make inferences about the market in the right panel.
@@
Limit FPPE. We first introduce our notion of a limit market and two regularity conditions on the market, which models the underlying market structure that we sample from. We have $n$ buyers and a possibly continuous set of items $\Theta$ with an integrating measure $\mathop{}\!\mathrm{d}\theta$. For example, one could take $\Theta=[0,1]$ and $\mathop{}\!\mathrm{d}\theta=$ the Lebesgue measure on $[0,1]$. Each buyer has a budget $b_{i}$; let $\bm{b}=(b_{1},\dots,b_{n})$. The valuation for buyer $i$ is a function $v_{i}\in L^{1}_{+}$ such that buyer $i$ has valuation $v_{i}(\theta)$ for one unit of item $\theta\in\Theta$; let $\bm{v}:\Theta\to{\mathbb{R}^{n}}$, $\bm{v}(\theta)=[v_{1}(\theta),\dots,v_{n}(\theta)]$. We assume ${\bar{v}}=\max_{i}\sup_{\theta}v_{i}(\theta)<\infty$. The supplies of items are given by a function $s\in L^{\infty}_{+}$, i.e., item $\theta\in\Theta$ has $s(\theta)$ units of supply. Without loss of generality, we assume a unit total supply $\int_{\Theta}s\mathop{}\!\mathrm{d}\theta=1$. Given $g:\Theta\to{\mathbb{R}}$, we let ${\mathbb{E}}[g]=\int g(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta$ and ${{Var}}[g]={\mathbb{E}}[g^{2}]-({\mathbb{E}}[g])^{2}$. Given $t$ i.i.d. draws $\{\theta^{1},\dots,\theta^{t}\}$ from $s$, let $P_{t}g({\cdot})=\frac{1}{t}{\sum_{\tau=1}^{t}}g({\theta^{\tau}})$.
@@
Next we introduce the market equilibrium concept that is the foundation of our study. For that we leverage the first-price pacing equilibrium (FPPE) (Conitzer et al., 2022a). FPPE models equilibrium outcomes under budget-management systems employed in several practical settings. Each buyer is assigned a pacing multiplier $\beta_{i}$, which is used to control their budget expenditure. For each individual auction $\theta$, the buyer bids $\beta_{i}v_{i}(\theta)$, which can be seen as their adjusted valuation after factoring in their budget constraint (in practice, the valuation $v_{i}(\theta)$ may not be the buyer’s true valuation, but instead their unscaled bid reported to the platform). The goal of the budget management system is to achieve an equilibrium, meaning that it must ensure that buyers spend their budget exactly by appropriately choosing $\beta_{i}$. If the budget cannot be fully spent, then no pacing must occur (i.e., $\beta_{i}=1$). Below we formally define the pacing equilibrium concept in the continuous setting (see Fig 2, right), and give the finite version in the next section (Fig 2, left).
Definition 1 (Limit FPPE, Gao & Kroer (2022); Conitzer et al. (2022a)).
@@
A limit FPPE, denoted as ${\mathsf{{FPPE}}}(\bm{b},\bm{v},s,\Theta)$, is the unique tuple $({\bm{\beta}},p({\cdot}))\in[0,1]^{n}\times L^{1}_{+}(\Theta)$ such that there exist an allocation $x_{i}:\Theta\to[0,1]$ for all $i\in[n]$ such that,
  1. 1.
@@
    (First-price) Prices and allocations are determined by first price auctions: for all items $\theta\in\Theta$, $p(\theta)=\max_{i}{\beta_{i}}v_{i}(\theta)$, and only the highest bidders obtain items: for all $i$ and $\theta$, $x_{i}(\theta)>0$ implies ${\beta_{i}}{v_{i}(\theta)}=\max_{k}\beta_{k}v_{k}(\theta)$
  2. 2.
@@
    (Feasiblity, market clearing) Budget are respected: for all $i$, $\int x_{i}(\theta)p(\theta)s(\theta)\mathop{}\!\mathrm{d}\theta\leq b_{i}$. There is no overselling: for all $\theta$, ${\sum_{i=1}^{n}}x_{i}(\theta)\leq 1$. Items with nonzero price are fully allocated: for all $\theta$, $p(\theta)>0$ implies ${\sum_{i=1}^{n}}x_{i}(\theta)=1$.
  3. 3.